1) True or False? A researcher applies a simple regression to get the results shown below using n=8 observations. Then, to construct the 95 percent confidence interval for the slope, we must use a t statistic of 2.447, by Appendix D.

Variable

Coefficients

Standard Error

Intercept

-0.1667

2.8912

X Variable (slope)

1.8333

0.2307

2) Based off the table presented above,  A researcher applies a simple regression to get the results shown below using n=8 observations. Which of the followings is the 95 percent confidence interval for the estimated slope?

a) [1.333,2.284]

b) [1.269,2.398]

c) [1.602,2.064]

d) [1.118,2.449]
Answer #1

a) test statistic for slope

T =( b1- B1 )/SEb1= 1.833/0.2307= 7.95

False.

b)Alpha=0.05 , n=8 , df= n-2 = 6

95 percent confidence interval for the estimated slope

Beta 1 +/- [ t crit * SEBeta1]

1.833 +/- [ 2.4469 * 0.2307 ]

1.833 +/- 0.5645

1.269 < beta < 2.398

Answer :- [ B ]